# GHZ protocols enhance frequency metrology despite spontaneous decay

### v1

@@ Abstract
We have presented a protocol with low complexity that saturates the QCRB of GHZ states and thus, unexpectedly, results in a significant enhancement compared to the SQL in the presence of spontaneous decay. This is achieved by a measurement and estimation scheme that allows to identify and exclude spontaneous emission events in the Ramsey sequence and thus implements an error detection and mitigation scheme to improve frequency metrology. Furthermore, the robustness was shown through comprehensive Monte-Carlo simulations of atomic clocks, thereby paving the way for near-term implementations into experimental setups. GHZ states are an attractive experimental choice, since in a variation of the protocol involving the entanglement of Zeeman states of equal magnitude but opposite sign, they can be made first order magnetic field insensitive [80], thus eliminate dephasing, the major source of decoherence [47]. Furthermore, the shorter optimum probe time compared to CSS reduces loss of contrast and time dilation shifts from motional heating of the ion crystal, thus improving the signal-to-noise ratio and the accuracy of such a clock. An open question remains to identify the optimal protocols saturating the OQI. However, evidence from numerical optimization suggests that large circuit depths may be required, similar to Refs. [48, 49]. In this context, investigating the transition from GHZ-like protocols for small ensembles to SSS in the asymptotic limit is of fundamental importance. In addition to frequency metrology, the findings of this work can be applied to various precision measurements where phase estimation is limited by spontaneous decay.
@@ Methods
$P\left(x|\phi,T\right)=\frac{1}{4}\binom{N}{N_{-}}\left[e^{\Gamma T(N-N_{-})}\left(1-e^{-\Gamma T}\right)^{N_{-}}\right.\\ +\left.e^{\Gamma TN_{-}}\left(1-e^{-\Gamma T}\right)^{N-N_{-}}\right]$ (16)
@@
and thus the QCRB is not saturated for $N>2$. Nevertheless, an improvement compared to the SQL is obtained which, however, vanishes in the asymptotic limit (cf. Fig. 2).
@@
In Sec. S1 the SQL (2) in the presence of dephasing and spontaneous decay is derived as well as the ultimate lower bound given by Eq. (6). Properties of the GHZ protocols are obtained in Sec. S2, namely the evaluation of the QCRB (3), the sensitivities of the parity-GHZ (14), linear-GHZ (17), as well as heralded-GHZ scheme, including Eqs. (4), (5), (15) and (16). In Sec. S3 the asymptotic scaling of the OQI is discussed.
@@
$+\left[e^{-\Gamma T}|\negmedspace\uparrow\rangle\langle\uparrow\negmedspace|+\left(1-e^{-\Gamma T}\right)|\negmedspace\downarrow\rangle\langle\downarrow\negmedspace|\right]^{\otimes N}\Big{)}.$ (S25)
@@
$+e^{-\frac{\Gamma+\gamma}{2}NT}\Big{[}e^{iN\phi}(-i)^{N+E}\Big{(}|\negmedspace\downarrow\rangle\langle\uparrow\negmedspace|^{\otimes N}+(-i)^{N+E}|\negmedspace\uparrow\rangle\langle\uparrow\negmedspace|^{\otimes N}+i^{N+E}|\negmedspace\downarrow\rangle\langle\downarrow\negmedspace|^{\otimes N}+|\negmedspace\uparrow\rangle\langle\downarrow\negmedspace|^{\otimes N}\Big{)}$ (S28)
$+e^{-iN\phi}i^{N+E}\Big{(}|\negmedspace\uparrow\rangle\langle\downarrow\negmedspace|^{\otimes N}+(-i)^{N+E}|\negmedspace\downarrow\rangle\langle\downarrow\negmedspace|^{\otimes N}+i^{N+E}|\negmedspace\uparrow\rangle\langle\uparrow\negmedspace|^{\otimes N}+|\negmedspace\downarrow\rangle\langle\uparrow\negmedspace|^{\otimes N}\Big{)}\Big{]}$ (S29)
$+\left[e^{-\Gamma T}|\negmedspace\uparrow\rangle\langle\uparrow\negmedspace|+\left(1-e^{-\Gamma T}\right)|\negmedspace\downarrow\rangle\langle\downarrow\negmedspace|\right]^{\otimes N}+i^{N+E}\left[e^{-\Gamma T}|\negmedspace\downarrow\rangle\langle\uparrow\negmedspace|+\left(1-e^{-\Gamma T}\right)|\negmedspace\uparrow\rangle\langle\downarrow\negmedspace|\right]^{\otimes N}$ (S30)
$+(-i)^{N+E}\left[e^{-\Gamma T}|\negmedspace\uparrow\rangle\langle\downarrow\negmedspace|+\left(1-e^{-\Gamma T}\right)|\negmedspace\downarrow\rangle\langle\uparrow\negmedspace|\right]^{\otimes N}+\left[e^{-\Gamma T}|\negmedspace\downarrow\rangle\langle\downarrow\negmedspace|+\left(1-e^{-\Gamma T}\right)|\negmedspace\uparrow\rangle\langle\uparrow\negmedspace|\right]^{\otimes N}\Big{\}}.$ (S31)
On $\rho_{\mathrm{final}}$ a projective measurement of $S_{z}$ is performed. We will refer to this interrogation scheme combined with the standard linear estimator as ’linear-GHZ’ protocol. The signal reads
@@
and thus yields a lower phase estimation uncertainty than the SQL. For $N=2$ the linear-GHZ protocol saturates the QCRB. However, for $N>2$ the QCRB is not saturated and the SQL is asymptotically approximated (cf. Fig. 2).
@@
$P\left(x|\phi,T\right)=$$\begin{cases}\frac{1}{4}\left[1+\left(1-e^{-\Gamma T}\right)^{N}+e^{-\Gamma NT}\mp 2e^{-\frac{\Gamma+\gamma}{2}NT}\cos(N\phi)\right]&\text{if }x=\pm\frac{N}{2}\\ \frac{1}{4}\binom{N}{N_{-}}\left[e^{\Gamma T(N-N_{-})}\left(1-e^{-\Gamma T}\right)^{N_{-}}+e^{\Gamma TN_{-}}\left(1-e^{-\Gamma T}\right)^{N-N_{-}}\right]&\text{if }x=\frac{N}{2}-N_{-}\\ \end{cases}$ (S35)
@@
$(\Delta\phi(T))^{2}=$ $\frac{N^{2}}{4}\frac{P\left(x=+\frac{N}{2}|\phi_{0},T\right)+P\left(x=-\frac{N}{2}|\phi_{0},T\right)}{\partial_{\phi}\langle X(\phi,T)\rangle|_{\phi=\phi_{0}}}$ (S36)
@@
S3 Asymptotic OQI
@@
While no general expressions for the OQI exist in the presence of decoherence and numerical optimization becomes challenging with increasing ensemble size $N$, in the asymptotic limit ($N\gg 1$) lower bounds on the sensitivity of the OQI can be derived for particular decoherence processes. In Ref. [44] a lower bound for individual dephasing was derived that yields a maximal improvement of $1/e$ over the SQL, considering arbitrary input states and a projective measurement of the spin in a suitable direction. The extension to spontaneous decay results in the same improvement, and the lower limit is given by Eq. (6), derived in Sec. S1. Interestingly, this already represents the asymptotic ultimate lower bound as derived in Refs. [43, 42, 76, 77] where the tightness was additionally proved. Generalizing the investigations of Ref. [41] to additionally include spontaneous decay (cf. Sec. S1), Eq. (6) is asymptotically saturated by squeezed spin states (SSS), e.g. generated by OAT interactions. Note that in the presence of spontaneous emission a lower bound smaller by a factor of $4$ compared to Eq. (6) was derived in Refs. [42, 76] which might be achievable by means of additional ancilla systems or adaptive quantum feedback strategies as quantum error correction schemes. However, saturability remains an open question and is beyond the scope of this work. Refs. [45, 75] give a comprehensive overview over different quantum control strategies and corresponding bounds.

### v2

@@ Abstract
We have presented a protocol with low complexity that saturates the QCRB of GHZ states and thus, unexpectedly, results in a substantial enhancement of 1.8 dB compared to the SQL in the presence of spontaneous decay. This is achieved by a measurement and estimation scheme that allows to identify and exclude spontaneous emission events in the Ramsey sequence and thus implements an error detection and mitigation scheme to improve frequency metrology. Additionally, we have identified a GHZ-like protocol saturating the OQI for small ensembles and performs closely to it for intermediate ensemble sizes. The 2.25 dB gain compared to the SQL results from an unequal superposition of the two maximal Dicke states, which counteracts spontaneous decay during the free evolution time and can be generated by a minor modification of the GHZ state. Furthermore, the robustness of these protocols was shown through comprehensive Monte-Carlo simulations of atomic clocks, thereby paving the way for near-term implementations into experimental setups. GHZ(-like) states are an attractive experimental choice, since in a variation of the protocol involving the entanglement of Zeeman states of equal magnitude but opposite sign, they can be made first order magnetic field insensitive [80], thus eliminate dephasing, the major source of decoherence [46]. Furthermore, the shorter optimum probe time compared to CSS reduces loss of contrast and time dilation shifts from motional heating of the ion crystal, thus improving the signal-to-noise ratio and the accuracy of such a clock. An open question remains to investigate the transition from GHZ-like protocols for small ensembles to SSS in the asymptotic limit. In addition to frequency metrology, the findings of this work can be applied to various precision measurements where phase estimation is limited by spontaneous decay.
@@ Methods
$P\left(x|\phi,T\right)=\frac{1}{4}\binom{N}{N_{-}}\left[e^{-\Gamma T(N-N_{-})}\left(1-e^{-\Gamma T}\right)^{N_{-}}\right.\\ +\left.e^{-\Gamma TN_{-}}\left(1-e^{-\Gamma T}\right)^{N-N_{-}}\right]$ (18)
@@
and thus the QCRB is not saturated for $N>2$. Nevertheless, an improvement compared to the SQL is obtained which, however, vanishes in the asymptotic limit (cf. Fig. 2B).
@@
In Sec. S1 the SQL (2) in the presence of dephasing and spontaneous decay is derived as well as the ultimate lower bound given by Eq. (6). Properties of the GHZ protocols are obtained in Sec. S2, namely the evaluation of the QCRB (3), the sensitivities of the parity-GHZ (16), linear-GHZ (19), heralded-GHZ as well as heralded-uGHZ scheme (8), including Eqs. (4), (5), (7), (17) and (18). The notion of quantum jumps is briefly presented in Sec. S3. In Sec. S4 the OQI and its asymptotic scaling is discussed.
@@
$\quad+\left[e^{-\Gamma T}|\negmedspace\uparrow\rangle\langle\uparrow\negmedspace|+\left(1-e^{-\Gamma T}\right)|\negmedspace\downarrow\rangle\langle\downarrow\negmedspace|\right]^{\otimes N}\Big{)}.$ (S25)
@@
$\quad+e^{-\frac{\Gamma+\gamma}{2}NT}\Big{[}e^{iN\phi}(-i)^{N+E}\Big{(}|\negmedspace\downarrow\rangle\langle\uparrow\negmedspace|^{\otimes N}+(-i)^{N+E}|\negmedspace\uparrow\rangle\langle\uparrow\negmedspace|^{\otimes N}+i^{N+E}|\negmedspace\downarrow\rangle\langle\downarrow\negmedspace|^{\otimes N}+|\negmedspace\uparrow\rangle\langle\downarrow\negmedspace|^{\otimes N}\Big{)}$ (S28)
$\quad+e^{-iN\phi}i^{N+E}\Big{(}|\negmedspace\uparrow\rangle\langle\downarrow\negmedspace|^{\otimes N}+(-i)^{N+E}|\negmedspace\downarrow\rangle\langle\downarrow\negmedspace|^{\otimes N}+i^{N+E}|\negmedspace\uparrow\rangle\langle\uparrow\negmedspace|^{\otimes N}+|\negmedspace\downarrow\rangle\langle\uparrow\negmedspace|^{\otimes N}\Big{)}\Big{]}$ (S29)
$\quad+\left[e^{-\Gamma T}|\negmedspace\uparrow\rangle\langle\uparrow\negmedspace|+\left(1-e^{-\Gamma T}\right)|\negmedspace\downarrow\rangle\langle\downarrow\negmedspace|\right]^{\otimes N}+i^{N+E}\left[e^{-\Gamma T}|\negmedspace\downarrow\rangle\langle\uparrow\negmedspace|+\left(1-e^{-\Gamma T}\right)|\negmedspace\uparrow\rangle\langle\downarrow\negmedspace|\right]^{\otimes N}$ (S30)
$\quad+(-i)^{N+E}\left[e^{-\Gamma T}|\negmedspace\uparrow\rangle\langle\downarrow\negmedspace|+\left(1-e^{-\Gamma T}\right)|\negmedspace\downarrow\rangle\langle\uparrow\negmedspace|\right]^{\otimes N}+\left[e^{-\Gamma T}|\negmedspace\downarrow\rangle\langle\downarrow\negmedspace|+\left(1-e^{-\Gamma T}\right)|\negmedspace\uparrow\rangle\langle\uparrow\negmedspace|\right]^{\otimes N}\Big{\}}.$ (S31)
On $\rho_{\mathrm{final}}$ a projective measurement of $S_{z}$ is performed. We will refer to this interrogation scheme combined with the standard linear estimator as ‘linear-GHZ’ protocol. The signal reads
@@
and thus yields a lower phase estimation uncertainty than the SQL. For $N=2$ the linear-GHZ protocol saturates the QCRB. However, for $N>2$ the QCRB is not saturated and the SQL is asymptotically approximated (cf. Fig. 2B).
@@
$P\left(x|\phi,T\right)=\begin{cases}\frac{1}{4}\left[1+\left(1-e^{-\Gamma T}\right)^{N}+e^{-\Gamma NT}\mp 2e^{-\frac{\Gamma+\gamma}{2}NT}\cos(N\phi)\right]&\text{if }x=\pm\frac{N}{2}\\ \frac{1}{4}\binom{N}{N_{-}}\left[e^{-\Gamma T(N-N_{-})}\left(1-e^{-\Gamma T}\right)^{N_{-}}+e^{-\Gamma TN_{-}}\left(1-e^{-\Gamma T}\right)^{N-N_{-}}\right]&\text{if }x=\frac{N}{2}-N_{-}\\ \end{cases}$ (S35)
@@
$(\Delta\phi(T))^{2}=\frac{N^{2}}{4}\frac{P\left(x=+\frac{N}{2}|\phi_{0},T\right)+P\left(x=-\frac{N}{2}|\phi_{0},T\right)}{\left(\partial_{\phi}\langle X(\phi,T)\rangle|_{\phi=\phi_{0}}\right)^{2}}$ (S36)
@@
Heralded-uGHZ protocol —
The initial state (7) is obtained by applying the rotation $\mathcal{R}_{z}(\theta)$ and OAT interaction $\mathcal{U}_{\mathrm{GHZ}}$ to the state (S24)
$|\psi_{\mathrm{in}}\rangle=\mathcal{U}_{\mathrm{GHZ}}\mathcal{R}_{z}(\theta)\mathcal{U}_{\mathrm{GHZ}}|\negmedspace\downarrow\rangle^{\otimes N}=\alpha_{\theta}|\negmedspace\downarrow\rangle^{\otimes N}+\beta_{\theta}|\negmedspace\uparrow\rangle^{\otimes N}=e^{-i\frac{\pi}{2E}}\left[i\sin\left(\tfrac{\theta N}{2}\right)|\negmedspace\downarrow\rangle^{\otimes N}+i^{N+E}\cos\left(\tfrac{\theta N}{2}\right)|\negmedspace\uparrow\rangle^{\otimes N}\right].$ (S37)
Fortunately, the remaining evaluation is analogous to the heralded-GHZ protocol but with modified coefficients $\alpha_{\theta}$ and $\beta_{\theta}$ of the maximal Dicke states $|\negmedspace\downarrow\rangle^{\otimes N}$ and $|\negmedspace\uparrow\rangle^{\otimes N}$. The optimal rotation angle reads
$\theta=\frac{2}{N}\arctan\left(\sqrt[4]{e^{-N\Gamma T}+\left(1-e^{-\Gamma T}\right)^{N}}\right)$ (S38)
and the associated minimal phase estimation uncertainty is given by (8), i.e. $(\Delta\phi(T))^{2}=\frac{e^{(\Gamma+\gamma)NT}}{4N^{2}}\left[1+\sqrt{e^{-N\Gamma T}+\left(1-e^{-\Gamma T}\right)^{N}}\right]^{2}$. Moreover, the corresponding initial state (S37) has populations
$P_{\mathrm{in}}\left(-\frac{N}{2}\right)$ $=\mathrm{Tr}\left(|\negmedspace\downarrow\rangle\langle\downarrow\negmedspace|^{\otimes N}|\psi_{\mathrm{in}}\rangle\langle\psi_{\mathrm{in}}|\right)=\frac{\sqrt{e^{-N\Gamma T}+\left(1-e^{-\Gamma T}\right)^{N}}}{1+\sqrt{e^{-N\Gamma T}+\left(1-e^{-\Gamma T}\right)^{N}}}$ (S39)
$P_{\mathrm{in}}\left(+\frac{N}{2}\right)$ $=\mathrm{Tr}\left(|\negmedspace\uparrow\rangle\langle\uparrow\negmedspace|^{\otimes N}|\psi_{\mathrm{in}}\rangle\langle\psi_{\mathrm{in}}|\right)=\frac{1}{1+\sqrt{e^{-N\Gamma T}+\left(1-e^{-\Gamma T}\right)^{N}}}$ (S40)
and consequently indeed represents a non-equal superposition, with a higher population in the excited state. Nevertheless, again the phase is only encoded in the maximal Dicke states after the free evolution time and the final OAT interaction
$P\left(x|\phi,T\right)=\begin{cases}\frac{1}{2}\sqrt{e^{-N\Gamma T}+\left(1-e^{-\Gamma T}\right)^{N}}\pm e^{-\frac{\Gamma+\gamma}{2}NT}\frac{\sqrt[4]{e^{-N\Gamma T}+\left(1-e^{-\Gamma T}\right)^{N}}}{1+\sqrt{e^{-N\Gamma T}+\left(1-e^{-\Gamma T}\right)^{N}}}\sin(N\phi)&\text{if }x=\pm\frac{N}{2}\\ \frac{1}{2}\frac{1}{1+\sqrt{e^{-N\Gamma T}+\left(1-e^{-\Gamma T}\right)^{N}}}\binom{N}{N_{-}}\left[e^{-\Gamma T(N-N_{-})}\left(1-e^{-\Gamma T}\right)^{N_{-}}+e^{-\Gamma TN_{-}}\left(1-e^{-\Gamma T}\right)^{N-N_{-}}\right]&\text{if }x=\frac{N}{2}-N_{-}\\ \end{cases}$ (S41)
and thus suggests the use of the dedicated estimator (5).
Note that the optimal measurement and estimation schemes for the heralded-GHZ and heralded-uGHZ protocol were discovered by a systematic study of a variational class, similar to Refs. [47, 48].
S3 Quantum jumps
In the following we will discuss the notion of quantum jumps in open quantum systems interacting with their environment based on master equations. In particular, we aim to identify and explain the ‘no-jump dynamics’ which is crucial for the presented GHZ protocols. The dynamics based on a general master equation is governed by
$\dot{\rho}=-i[H,\rho]+\left(A\rho A^{\dagger}-\frac{1}{2}A^{\dagger}A\rho-\frac{1}{2}\rho A^{\dagger}A\right)$ (S42)
with Hamiltonian $H$ and Lindblad (jump) operator $A$ representing an arbitrary dissipative process (quantum jump). For simplicity, we will consider a single jump operator, although the method can easily be generalized to include several noise processes. In our framework, the Hamiltonian $H=\omega S_{z}$ is given by the frequency deviation $\omega$ and spin operator along the $z$-direction $S_{z}$, while the jump operators for each qubit are $A=\sqrt{\Gamma}\sigma_{-}$ and $A=\frac{\sqrt{\gamma}}{2}\sigma_{z}$ for spontaneous decay and individual dephasing, respectively. It is instructive to rewrite the master equation according to
$\dot{\rho}=-i(H_{\mathrm{eff}}\rho-\rho H_{\mathrm{eff}}^{\dagger})+A\rho A^{\dagger}\qquad\text{with}\qquad H_{\mathrm{eff}}=H-\frac{i}{2}A^{\dagger}A.$ (S43)
The dynamics can be interpreted as a combination of a continuous and a stochastic evolution. The continuous evolution, also called ‘no-jump dynamics’, is determined by the effective Hamiltonian $H_{\mathrm{eff}}$ and does not imply any quantum jumps, i.e. no quanta are exchanged with the environment. Nevertheless, an effective decay is caused by the non-hermitian part of $H_{\mathrm{eff}}$. In contrast, the stochastic contribution is characterized by the term $A\rho A^{\dagger}$ and results in sudden quantum jumps, interrupting the continuous evolution. These jumps reflect discrete events like the emission of a photon.
A series expansion of the formal solution to the the master equation (S43) is given by
$\rho(t,t_{0})=G(t,t_{0})\rho(t_{0})+\int_{t_{0}}^{t}\mathrm{d}t_{1}G(t,t_{1})JG(t_{1},t_{0})\rho(t_{0})+\int_{t_{0}}^{t}\mathrm{d}t_{2}\int_{t_{0}}^{t}\mathrm{d}t_{1}G(t,t_{2})JG(t_{2},t_{1})JG(t_{1},t_{0})\rho(t_{0})+\ldots$ (S44)
with superoperators $G(t,t_{0})B=e^{-iH_{\mathrm{eff}}(t-t_{0})}Be^{iH_{\mathrm{eff}}^{\dagger}(t-t_{0})}$ and $JB=ABA^{\dagger}$ representing the continuous non-unitary time propagation and the quantum jump, respectively. The terms in Eq. (S44) include an increasing number of jumps (characterized by the number of $J$ operators) in the time interval $[t_{0},t]$. In particular, the first term $G(t,t_{0})\rho(t_{0})$ does not include a single jump and thus is referred to as ‘no-jump dynamics’.
S4 Optimal quantum interferometer
The optimal quantum interferometer (OQI) defines the ultimate lower limit in phase estimation for a particular setup. Considering a phase evolution implemented by an unitary rotation $\mathcal{R}_{z}(\phi)$, the OQI $(\Delta\phi_{\mathrm{OQI}}(T))^{2}=1/\mathcal{F}_{\mathrm{OQI}}$ can be determined by optimizing the QFI over all input states
$\mathcal{F}_{\mathrm{OQI}}=\sup_{|\psi\rangle}\mathcal{F}_{Q}\left[\Lambda_{T}[|\psi\rangle\langle\psi|],S_{z}\right]$ (S45)
where $\Lambda_{T}$ describes any decoherence processes. In an ideal scenario, i.e. $\Lambda_{T}[|\psi\rangle\langle\psi|]=|\psi\rangle\langle\psi|$, the well known Heisenberg limit $\mathcal{F}_{\mathrm{OQI}}=N^{2}$ is obtained. However, no general expressions for the OQI exist in the presence of decoherence and a numerical evaluation is necessary.
Numerical computation —
Considering permutational invariance of the ensemble, a basis is given by the $n_{\mathrm{PI}}=\left(\frac{N}{2}+1\right)^{2}-\frac{1}{4}\mathrm{mod}_{2}(N)$ permutationally invariant Dicke states, resulting in a considerably smaller subspace $\mathcal{O}(N^{2})$ compared to the full $2^{N}$-dimensional Hilbert space. Additionally assuming a pure initial state with total spin $S=N/2$, $|\psi\rangle$ contains $N+1$ non-zero coefficients at most. Nevertheless, a brute force optimization is very inefficient. A more efficient approach to compute the OQI is presented in Ref. [58], which iteratively optimizes the initial probe state and the observable. Starting from an arbitrary state, repeatedly the optimal measurement (via the QFI) and the corresponding optimal probe state are determined. Both involve the time evolution with $\Lambda_{T}$ and a matrix diagonalization. Fortunately, the dynamics for a fixed interrogation time is identical in each step and thus can be evaluated once in advance. Accordingly, the computational complexity of the time evolution in each step can be neglected. Since the permutationally invariant subspace has dimension $\mathcal{O}(N^{2})$, matrix diagonalization of the state and observable has complexity $\mathcal{O}(N^{6})$. Note, however, that in practice, the complexity might be reduced due to the block-diagonal structure of the states and spin operators, resulting in only $\mathcal{O}(N^{3})$ of the $\mathcal{O}(N^{4})$ elements being non-zero. Hence, the complexity of a single step in the iterative procedure is $\mathcal{O}(N^{6})$ at most. Consequently, despite its polynomial scaling, the iterative algorithm becomes challenging in terms of computational time with increasing ensemble size $N$. Additionally, the choice of the first state in the algorithm might affect convergence and result in local extrema. Moreover, numerical errors increase with $N$. Altogether, the evaluation of the OQI and minimization with respect to the interrogation time is only feasible for a few tens of qubits on a standard PC in the presence of decoherence.
Asymptotic scaling —
@@
In contrast, in the asymptotic limit ($N\gg 1$) lower bounds on the sensitivity of the OQI can be derived for particular decoherence processes. In Ref. [43] a lower bound for individual dephasing was derived that yields a maximal improvement of $1/e$ over the SQL, considering arbitrary input states and a projective measurement of the spin in a suitable direction. The extension to spontaneous decay results in the same improvement, and the lower limit is given by Eq. (6), derived in Sec. S1. Interestingly, this already represents the asymptotic ultimate lower bound as derived in Refs. [42, 41, 76, 77] where the tightness was additionally proved. Generalizing the investigations of Ref. [40] to additionally include spontaneous decay (cf. Sec. S1), Eq. (6) is asymptotically saturated by squeezed spin states (SSS), e.g. generated by OAT interactions. Note that in the presence of spontaneous emission a lower bound smaller by a factor of $4$ compared to Eq. (6) was derived in Refs. [41, 76] which might be achievable by means of additional ancilla systems or adaptive quantum feedback strategies as quantum error correction schemes. However, saturability remains an open question and is beyond the scope of this work. Refs. [44, 75] give a comprehensive overview over different quantum control strategies and corresponding bounds.